OpenStudy (anonymous):

solve. 25^(z - 4) = 125

OpenStudy (anonymous):

not one of my choices

OpenStudy (anonymous):

i got -2.5

OpenStudy (anonymous):

Start by taking the natural log of both sides. This will drop the exponent down so you can solve for z: \[(z-4) \ln25 = \ln 125\]How can you solve for z now?

OpenStudy (anonymous):

ad 4 to both sides

OpenStudy (anonymous):

Not yet, first we have to get rid of the ln 25 on the left side. So we'll divide both sides by ln 25: \[z-4=\frac{ \ln 125 }{ \ln 25 }\]

OpenStudy (anonymous):

ooooh ok i see ya!